# DynST: Dynamic Sparse Training for Resource-Constrained Spatio-Temporal Forecasting

### v1

@@ 1. Introduction
Traditional approaches to sensor running time optimization (Priyadarshi et al., 2020; Zou and Chakrabarty, 2003; Yarinezhad and Hashemi, 2023; Xu, 2020; Kundu and Das, 2023), e.g., virtual force and Voronoi diagrams, utilize specific algorithms to dynamically activate sensors. These methods dynamically adjust the activation times of sensors to optimize the detection process across each sub-region. Unfortunately, generating an effective activation strategy using only pre-existing historical observation data or urban geographic characteristics is very tricky, as it often involves complex technical design (Zhang et al., 2023a). Furthermore, with numerous factors influencing sensor deployment, relying solely on single variables (such as urban layout or geographic features) does not accurately capture the optimal deployment strategy (Yan and Li, 2023; Zheng et al., 2023).
With this in mind, in this paper, our aim is to speedup inference time by proposing a novel sensor deactivation strategy, which is based on historical observations. A promising direction involves adopting deep-learning-oriented metrics to adaptively and dynamically evaluate or verify the benefits brought by each sensor deployment. The ever-increasing dynamic sparse training (termed DST) (Evci et al., 2020; Liu et al., 2021; Huang et al., 2023; Liu et al., 2020), though opening a potential path for the upcoming automating effective deployment, sadly drops a daunting obstacle on the way towards their spatio-temporal on-device deployment. Concretely, DST technology demonstrates the potential to train a sub-network from scratch, using sparse network training, to match the performance of a fully dense network. In real world, the training of models and the optimization of sensors are still heavily in both academia and industry. Transferring the concept of DST to spatio-temporal forecasting realm is intuitively beneficial, as it can significantly accelerate model training while optimizing deployment.
@@
To bridge the gap between industry and academia, this paper introduces for the first time the concept of dynamic sparse training for spatio-temporal data, termed DynST. DynST dynamically trains to filter out the crucial parts of data for future predictions, and eliminates non-essential services to achieve resource-constrained service management. Concretely, DynST employs dynamic training to apply masking to historical regions, with the aim of aggressively reducing the proliferation of sensor deployment. This approach is taken at the algorithmic level to more effectively mask individual regions (each corresponding to a sensor device). Given the dynamic nature of time-series data, we utilize explicit channel stacking to construct overlapping saliency maps of historical regions. This facilitates the scoring of the importance of sensors in each region.
DynST is both simple and efficient, demonstrating powerful optimization capabilities across a variety of industrial scenarios. It effectively reduces historically insignificant observation areas (i.e., sub-regions) in both regular and inherently irregular data environments, without impacting the performance of future predictions.
@@
  • Significantly Improves Inference Efficiency. DynST enhances inference speed across different architectures. On the Turbulence dataset, the STGCN architecture speeds up by 72% to 1.721 times with DynST. On the Fire dataset, the GNN architecture speeds up by about 14.5% to 1.541 times. On the JD Taxibj+ dataset, the Transform architecture nearly doubles in speed, increasing by about 34.5% to 1.987 times. These examples demonstrate DynST’s ability to improve computational efficiency, speeding up inference and handling large datasets efficiently.
@@ 2. Related Work
ST predictive learning can be categorized into three main types. Convolutional Neural Network (CNN)-based architectures: This research focuses on spatial feature extraction using CNN-based structures (Gao et al., 2022b; Tan et al., 2022; Wu et al., 2023c; Shi et al., 2015). These architectures use convolutional layers to effectively detect patterns in image and video data. Key advancements include deep convolutional networks for complex feature extraction and 3D convolutions for spatial-temporal analysis in video processing (Wang et al., 2018b); Recurrent Neural Network (RNN)-based Architectures: RNNs are used to optimize temporal data handling (Wang et al., 2017, 2018a, 2019), which are key for tasks like sequence prediction and time-dependent data analysis; Transformer-based Architectures delve into Transformer-based architectures for spatio-temporal data handling (Gao et al., 2022a; Bai et al., 2022; Wu et al., 2023b, c), by employing their self-attention mechanism to effectively manage sequence data. They capture long-range dependencies in both spatial and temporal dimensions, making them suitable for complex sequence modeling and analysis. Notably, there are models that leverage graph neural networks primarily for ST graph management (Ji et al., 2023; Shao et al., 2022; Li et al., 2017), we will discuss later.
@@ 4.2. Image-type Data Notations
For effective modeling in image-type data $\mathcal{X}$, we initially divide the total urban area into $p\times p$ sub-regions (patches), with each patch encompassing $(H/p,W/p)$ pixels. $H$ and $W$ is the height and the width of the input images. It is worth noting that the choice of $p$ should balance the trade-off between practicality and spatial granularity. In our implementation, we partition the entire urban area into small squares, each comprising $p\times p$ sensors, adhering to practicality requirements.
@@ 4.3. Problem Formulation
where $s_{g}$ is the sparsity, $||\cdot||_{0}$ counts the number of non-zero elements, and $\epsilon$ is the threshold for permissible performance difference. *** denotes the graph or image inputs and $\mathcal{R}$ represents the evaluation metrics.
@@ 5. Method
Fig 2 illustrate the overview of DynST framework. In Earth sciences, sensor deployment typically falls into two categories, i.e., image- and graph-type. Image-type deployment ensures that each area (termed ‘patch") is well covered by a sensor, while in graph-type deployment, the information from a node can be understood as being collected by a single sensor. To demonstrate the universal capabilities of DynST, we systematically consider both of these deployment types and perform a patchify operation on the images (Wu et al., 2023a). For graph data, since nodes can be defined as sensors, in this study, we do not perform any operations at the data input stage.
@@ 5.2. Iterative Pruning towards High Sparsity
With $M_{g}$ at hand, we proceed to train the models together with the fixed input graph and the graph mask, denoted as $f(M_{g}\odot{\tilde{\mathcal{G}}_{in}},\mathbf{\Theta})$, $f$ denotes the mapping function of the input ST model. with the objective function in Eq. 2, we aim to gradually find the sparse sub-graph towards better semantical preservation. One promising approach is to adopt one-shot pruning (Ma et al., 2021; Frankle et al., 2020), however, the sparse mask acquired through one-shot pruning is suboptimal. In fact, the assessment of each sensor necessitates iterative testing to ensure that the removal of a specific area does not significantly impact future predictions. To achieve our objectives, we employ an iterative pruning strategy (Chen et al., 2021) to gradually increase network sparsity. Assuming that each pruning iteration trims $p\%$ of the data parameters, after $\phi$ rounds of pruning, the remaining regions exhibit distinct advantages over the one-shot approach–that is–By iteratively pruning and retraining, the network can more effectively identify which parts are less important, as the remaining parameters have undergone $\phi$ rounds of repeated verification. Unlike previous iterative pruning literature, we alternately train the network and the mask $M_{g}$ to ensure that the mask can fully assimilate the effective information from the training process:
@@ 5.3. Dynamical Sparse Training
Then, at the begin of the round $\omega+1$, we continue to trian and adjust the mask for sending it to $\omega+1$ round pruning. We binarize the mask $\mathbf{M}_{g}^{(\omega+1)}$ after another $\Delta T$ iteration training. Without loss of generality, taking the semi-supervised node classification task as an example, our objective function can be expressed as follows:
@@ 5.4. Model Summary & Complexity Analysis
For image-type data, we transform each sub-region into a patch, which can also be understood as the concept of a “node”. Therefore, by training in a similar manner, we can identify the important sub-regions accordingly. DynST can enhance the inference speed of the model, which specifically depends on the predefined sparsity $s_{g}\%$. Typically, this results in an acceleration ratio of $1/s_{g}\%$. We summarize our prospective system and algorithm in Fig 4 and Appendix C, respectively.
@@ 4. $\mathcal{RQ}$4:
  • Main experiment. We conduct a comprehensive comparative analysis on various scientific datasets, covering meteorology, combustion science, traffic studies, and turbulence dynamics. The study encompasses both mainstream Graph Neural Network (GNN) architectures and non-GNN structures. In the appendix B, we detail the methods for data preprocessing, including how to convert raw data into graphical and image formats.
  • Multiple Training Strategies Experiments. We choose Weatherbench as the benchmark dataset, to evaluate the effectiveness of DynST when combining different training schemes. Specifically, in the training phase, we not only consider the impacts of parallel prediction and autoregressive iterative prediction but also introduce iterative pruning and one-shot pruning strategies. We focus on assessing the impacts of these strategies on model size, computational efficiency, and accuracy.
  • Ablation experiment. We carry out comprehensive ablation studies on the Jingdong Technology industry-level traffic dataset, Taxibj+, to validate the impact of various design choices on the practical implementation of our model. Through these experiments, we aim to deeply understand how the DynST concept affects data interpretability and the overall effectiveness.
@@ 6.1. Dataset & Backbones
6.2. Main experiments ($\mathcal{RQ}$1)
Obs 1.DynST has demonstrated that the removal of certain parts from the input data does not affect the model’s performance. As shown in Tab 1, We can easily observe the outcomes following the integration of the DynST concept into the model (+DynST). In the GNN architecture, the addition of DynST generally has a minimal impact on MAE. For example, on the WeatherBench ♣ and FIT $\varphi$ datasets, the MAE slightly increases from 4.35 to 4.37 and from 0.92 to 0.95, respectively. In non-GNN architectures, DynST usually maintains or reduces the MAE. For instance, in the ViT architecture on the Taxibj+ Outflow dataset, the MAE decreases from 3.67$\rightarrow$3.59. In particular, DynST generally significantly enhances the inference speed across various architectures. For example, in WeatherBench ♣, STGCN speeds up to 1.721 times, EGNN on FIT $\varphi$ to 1.541 times, and ViT on Taxibj+ Outflow to 1.987 times, effectively boosting the efficiency of inference.
Obs 2. DynST shows high efficiency in several scenarios. DynST also highly effective in improving the inference efficiency of various architectures. For example, on the WeatherBench ♣ dataset, the inference speed of STGCN increased by 23.7% with DynST (from the original speed to 1.721 times faster). Similarly, on the FIT $\varphi$ dataset, the EGNN architecture achieved a 14.5% speed increase with DynST (reaching 1.541 times faster). Moreover, on the Taxibj+ Outflow dataset, the inference speed of the ViT architecture almost doubled, specifically a 34.5% increase (rising to 1.987 times faster). These examples collectively show DynST’s capability to significantly enhance computational efficiency in various scenarios. The percentage-based speed improvements highlight its notable advantage in accelerating the inference of various ST architectures.
6.3. Deep insights ($\mathcal{RQ}$2 & $\mathcal{RQ}$3)
In this section, we conduct a more systematic study of DynST’s ability to accelerate inference. We select both graph and image-type data to observe model performance at various levels of sparsity. Concretely, for graph-type data, we choose Taxibj+ and EAGLE as benchmarks. For image-type data, we choose temperature ($\phi$) variable of FIT datasets and the temperature (♣) variable of the WeatherBench as verification. We integrate it with existing general frameworks and set the iterative pruning process to occur 10 times, with each iteration reducing the data volume by $\{1\%,2\%,\cdots,6\%\}$. Then we can obtain the data sparsity $\{10\%,20\%,\cdots,60\%\}$. We employ roll out strategy (Luo et al., 2023) to iteratively predict long sequence and verify the long-term prediction ability of baselines after involve DynST. We list the observations as follow.
@@ 6.3. Deep insights ($\mathcal{RQ}$2 & $\mathcal{RQ}$3)
Figure 5. The performance visualization of FIT datasets.
Figure 6. The performance visualization of JD Technology Taxibj+ datasets.
Obs 4. DynST effectively meets industrial-level requirements (30% sparsity), helping to achieve manageable inference demands while reducing the burden of inference. As shown in Fig 5, The first line of the display meticulously captures the actual observed temperature flow field, providing a vivid and accurate representation of the existing conditions. In contrast, the second line offers a predictive perspective, showcasing the temperature flow field as forecasted by the innovative Earthfarseer+DynST model. This juxtaposition not only illustrates the capabilities of the predictive model but also allows for a direct comparison between observed and predicted states. Bottom: Delving deeper into the analysis, the left image opens a window into a detailed time series comparison. It meticulously charts both the real and the predicted temperatures at the specific coordinates of (50,7), offering a granular view of the model’s precision over time. Similarly, the right image extends this comparison to another set of coordinates, (425,7), revealing how the model captures the temporal evolution of temperatures in this distinct area. These results showcase the remarkable ability of the DynST-enhanced model to preserve high local fidelity. This fidelity is not just theoretical; it translates into practical, industry-level reliability, consistently maintaining the prediction deviation within a tight 15% margin (Verda et al., 2021). Such performance not only underscores the robustness of the Earthfarseer+DynST model but also highlights its potential for widespread application in scenarios demanding high precision and reliability (Fig 6 also support our research findings).

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/3690624.3709391ccs: Applied computing Environmental sciences
@@ 1. Introduction
Traditional approaches to sensor deployment optimization (Priyadarshi et al., 2020; Zou and Chakrabarty, 2003; Yarinezhad and Hashemi, 2023; Xu, 2020; Kundu and Das, 2023), such as virtual force and Voronoi diagrams, utilize specific algorithms to select the regions where sensors should be deployed. These methods ideally adjust the activation of sensors to optimize the detection process across each sub-region.
Research Gap. Unfortunately, generating an effective activation strategy using only pre-existing historical observation data or urban geographic characteristics is very tricky, as it often involves complex technical design (Zhang et al., 2023a). Furthermore, with numerous factors influencing sensor deployment, relying solely on single variables (such as urban layout or geographic features) does not accurately capture the optimal deployment strategy (Yan and Li, 2023; Zheng et al., 2023).
With this in mind, in this paper, our aim is to speedup inference and training time by proposing a novel sensor deactivation strategy, which is based on historical observations. A promising direction and motivation involves adopting deep-learning-oriented metrics to adaptively and dynamically evaluate or verify the benefits brought by each sensor deployment. The ever-increasing dynamic sparse training (termed DST) (Evci et al., 2020; Liu et al., 2021; Huang et al., 2023; Liu et al., 2020), though opening a potential path for automating effective deployment, still remains in its nascent stages when exploring spatio-temporal scenarios.
However, transferring the concept of Dynamic Sparse Training to the realm of spatio-temporal forecasting is intuitively beneficial as it can significantly accelerate model training while optimizing deployment. Specifically, DST technology shows promise in training a sub-network from scratch, employing sparse network training strategies, to achieve the performance levels of a fully dense network. In practical terms, if sensor deployment nodes are considered as intuitive data distribution collection points, both the training of models and the optimization of sensors remain computationally intensive tasks in both academic and industrial settings.
@@
To bridge the gap between industry and academia, this paper introduces for the first time the concept of dynamic sparse training for spatio-temporal data, termed DynST. DynST dynamically trains to filter out the crucial parts of data for future predictions, and eliminates non-essential services to achieve resource-constrained service management. Concretely, DynST utilizes dynamic training through a differentiable mask applied to historical regions, aiming to significantly reduce the proliferation of sensor deployment. This approach is taken at the algorithmic level to more effectively mask individual regions (each corresponding to a sensor device). Given the dynamic nature of time-series data, we utilize explicit channel stacking to construct overlapping saliency maps of historical regions. This facilitates the scoring of the importance of sensors in each region. DynST is both model-agnostic and efficient, demonstrating powerful optimization capabilities across a variety of industrial scenarios. It effectively reduces historically insignificant observation areas (i.e., sub-regions) in both regular and inherently irregular data environments, without impacting the performance of future predictions.
@@
  • Significantly Improves Inference Efficiency. DynST enhances inference speed across different architectures. On the Turbulence dataset, the STGCN architecture speeds up by 72% to 1.721 times with DynST. On the Fire dataset, the GNN architecture speeds up by about 14.5% to 1.541 times. On the JD Taxibj+ dataset, the Transform architecture nearly doubles in speed, increasing by about 34.5% to 1.987 times. These examples demonstrate DynST’s ability to improve computational efficiency, speed up inference, and handle large datasets efficiently.
@@ 2. Related Work
ST predictive learning can be categorized into three main types. Convolutional Neural Network (CNN)-based architectures: This research focuses on spatial feature extraction using CNN-based structures (Gao et al., 2022b; Tan et al., 2022; Wu et al., 2023c; Shi et al., 2015). These architectures use convolutional layers to effectively detect patterns in image and video data. Key advancements include deep convolutional networks for complex feature extraction and 3D convolutions for spatial-temporal analysis in video processing (Wang et al., 2018b); Recurrent Neural Network (RNN)-based Architectures: RNNs are used to optimize temporal data handling (Wang et al., 2017, 2018a, 2019), which are key for tasks like sequence prediction and time-dependent data analysis; Transformer-based Architectures delve into Transformer-based architectures for spatio-temporal data handling (Gao et al., 2022a; Bai et al., 2022; Wu et al., 2023b, c), by employing their self-attention mechanism to effectively manage sequence data. They capture long-range dependencies in both spatial and temporal dimensions, making them suitable for complex sequence modeling and analysis. Notably, there are models that leverage graph neural networks primarily for ST graph management (Ji et al., 2023; Shao et al., 2022; Li et al., 2017), which we will discuss later.
@@ 4.2. Image-type Data Notations
For effective modeling in image-type data $\mathcal{X}$, we initially divide the total urban area into $p\times p$ sub-regions (patches), with each patch encompassing $(H/p,W/p)$ pixels. $H$ and $W$ are the height and the width of the input images. It is worth noting that the choice of $p$ should balance the trade-off between practicality and spatial granularity. In our implementation, we partition the entire urban area into small squares, each comprising $p\times p$ sensors, adhering to practicality requirements.
@@ 4.3. Problem Formulation
where $s_{g}$ is the sparsity, $||\cdot||_{0}$ counts the number of non-zero elements, and $\epsilon$ is the threshold for permissible performance difference. * denotes the graph or image inputs and $\mathcal{R}$ represents the evaluation metrics. The above equation expresses that under the sparsity level $s_{g}$, the model still maintains a loss comparable to that of predictions made with complete data, indicating that there is no significant performance degradation.
@@ 5. Method
Fig 2 illustrates the overview of the DynST framework. In Earth sciences, sensor deployment typically falls into two categories, i.e., image- and graph-type. Image-type deployment ensures that each area (termed ‘patch") is well covered by a sensor, while in graph-type deployment, the information from a node can be understood as being collected by a single sensor. To demonstrate the universal capabilities of DynST, we systematically consider both of these deployment types and perform a patchify operation on the images (Wu et al., 2023a). For graph data, since nodes can be defined as sensors, in this study, we do not perform any operations at the data input stage.
@@ 5.2. Iterative Pruning towards High Sparsity
With $M_{g}$ at hand, we proceed to train the models together with the fixed input graph and the graph mask, denoted as $f(M_{g}\odot{\tilde{\mathcal{G}}_{in}},\mathbf{\Theta})$, $f$ denotes the mapping function of the input ST model. with the objective function in Eq. 2, we aim to gradually find the sparse sub-graph towards better semantical preservation. One promising approach is to adopt one-shot pruning (Ma et al., 2021; Frankle et al., 2020). However, the sparse mask acquired through one-shot pruning is suboptimal. In fact, the assessment of each sensor necessitates iterative testing to ensure that the removal of a specific area does not significantly impact future predictions. To achieve our objectives, we employ an iterative pruning strategy (Chen et al., 2021) to gradually increase network sparsity. Assuming that each pruning iteration trims $p\%$ of the data parameters, after $\phi$ rounds of pruning, the remaining regions exhibit distinct advantages over the one-shot approach–that is–By iteratively pruning and retraining, the network can more effectively identify which parts are less important, as the remaining parameters have undergone $\phi$ rounds of repeated verification. Unlike previous iterative pruning literature, we alternately train the network and the mask $M_{g}$ to ensure that the mask can fully assimilate the effective information from the training process:
@@ 5.3. Dynamical Sparse Training
Then, at the beginning of the round $\omega+1$, we continue to train and adjust the mask for sending it to $\omega+1$ round pruning. We binarize the mask $\mathbf{M}_{g}^{(\omega+1)}$ after another $\Delta T$ iteration training. Without loss of generality, taking the semi-supervised node classification task as an example, our objective function can be expressed as follows:
@@ 5.4. Model Summary & Complexity Analysis
For image-type data, we transform each sub-region into a patch, which can also be understood as the concept of a “node”. Therefore, by training in a similar manner, we can identify the important sub-regions accordingly. DynST can enhance the inference speed of the model, which specifically depends on the predefined sparsity $s_{g}\%$. Typically, this results in an acceleration ratio of $1/s_{g}\%$. We summarize our algorithm in Alg.1.
Algorithm 1 Dynamic Sparse Training (DynST) Framework
0:  Input graph $\mathcal{G}_{in}$, Network $f$, Target Sparsity $S_{g}\%$
0:  Sparse mask $M_{g}$
1:  Initialize graph mask $M_{g}$
3:  while $1-\frac{\|M_{g}\|_{0}}{\|\mathcal{G}_{in}[;*]\|_{0}}<S_{g}$ do
4:     Training network for $R$ iterations
5:     Training $M_{g}$ for $M$ iterations
6:     Dynamical sparse training using Eq.5 and Eq.6
7:     Adjust $M_{g}$ using Eq.7
8:  end while
9:  $\hat{\mathcal{G}}_{in}\leftarrow\hat{\mathcal{G}}_{in}\odot M_{g}$
@@ 4. $\mathcal{RQ}$4:
  • Main experiment. We conduct a comprehensive comparative analysis on various scientific datasets, covering meteorology, combustion science, traffic studies, and turbulence dynamics. The study encompasses both mainstream Graph Neural Network (GNN) architectures and non-GNN structures (data pre-processing protocols are placed in appendix B).
  • Multiple Training Strategies Experiments. We choose Weatherbench as the benchmark, to evaluate the effectiveness of DynST when combining different training schemes. Specifically, in the training phase, we not only consider the impacts of parallel prediction and autoregressive iterative prediction but also introduce iterative pruning and one-shot pruning strategies.
  • Ablation experiment. We carry out comprehensive ablation studies on the Jingdong Technology industry-level traffic dataset, Taxibj+, to validate the impact of various design choices on the practical implementation of our model.
@@ 6.1. Dataset & Backbones
Figure 5. The performance visualization of FIT datasets. We can see that the overall temperature deviation is within 10 degrees Celsius, meeting the requirements of the fire science field. (Emmons, 1986)
Figure 6. The performance visualization of JD Technology Taxibj+ datasets. It displays the traffic inflow and outflow prediction results of various methods such as Earthfarseer, TAU, SimVP, and their versions combined with DnyST, along with their MAE scores. The left side shows the comparison of the Ground Truth with the predictions for each method, and the right side presents the outflow predictions.
6.2. Main experiments ($\mathcal{RQ}$1)
Obs 1.DynST has demonstrated that the removal of certain parts from the input data does not affect the model’s performance. As shown in Tab 1, We can easily observe the outcomes following the integration of the DynST concept into the model (+DynST). In the GNN architecture, the addition of DynST generally has a minimal impact on MAE. For example, on the WeatherBench ♣ and FIT $\varphi$ datasets, the MAE slightly increases from 4.35 to 4.37 and from 0.92 to 0.95, respectively. In non-GNN architectures, The addition of DynST still has a minimal impact on the MAE. For instance, in the ViT architecture on the Taxibj+ Outflow dataset, the MAE decreases from 3.67$\rightarrow$3.59. In particular, DynST generally significantly enhances the inference speed across various architectures. For example, in WeatherBench ♣, STGCN speeds up to 1.721 times, EGNN on FIT $\varphi$ to 1.541 times, and ViT on Taxibj+ Outflow to 1.987 times.
Obs 2. DynST shows high efficiency in several scenarios. DynST is also highly effective in improving the inference efficiency of various architectures. For example, on the WeatherBench ♣ dataset, the inference speed of STGCN increased by 23.7% with DynST (from the original speed to 1.721 times faster). Similarly, on the FIT $\varphi$ dataset, the EGNN architecture achieved a 14.5% speed increase with DynST (reaching 1.541 times faster). Moreover, on the Taxibj+ Outflow dataset, the inference speed of the ViT architecture almost doubled, specifically a 34.5% increase (rising to 1.987 times faster). These examples collectively show DynST’s capability to significantly enhance computational efficiency in various scenarios. The percentage-based speed improvements highlight its notable advantage in accelerating the inference of various ST architectures.
6.3. Deep insights ($\mathcal{RQ}$2 & $\mathcal{RQ}$3)
In this section, we conduct a more systematic study of DynST’s ability to accelerate inference. We select both graph and image-type data to observe model performance at various levels of sparsity. Concretely, for graph-type data, we choose Taxibj+ and EAGLE as benchmarks. For image-type data, we choose temperature ($\phi$) variable of FIT datasets and the temperature (♣) variable of the WeatherBench as verification. We integrate it with existing general frameworks and set the iterative pruning process to occur 10 times, with each iteration reducing the data volume by $\{1\%,2\%,\cdots,6\%\}$. Then we can obtain the data sparsity $\{10\%,20\%,\cdots,60\%\}$. We employ roll out strategy (Luo et al., 2023) to iteratively predict long sequence and verify the long-term prediction ability of baselines after involving DynST. We list the observations as follows.
@@ 6.3. Deep insights ($\mathcal{RQ}$2 & $\mathcal{RQ}$3)
Obs 4. DynST effectively meets industrial-level requirements (30% sparsity), helping to achieve manageable inference demands while reducing the burden of inference. As shown in Fig 5, The first line of the display meticulously captures the actual observed temperature flow field, providing a vivid and accurate representation of the existing conditions. In contrast, the second line offers a predictive perspective, showcasing the temperature flow field as forecasted by the innovative Earthfarseer+DynST model. This juxtaposition not only illustrates the capabilities of the predictive model but also allows for a direct comparison between observed and predicted states. Bottom: Delving deeper into the analysis, the left image opens a window into a detailed time series comparison. It meticulously charts both the real and the predicted temperatures at the specific coordinates of (50,7), offering a granular view of the model’s precision over time. Similarly, the right image extends this comparison to another set of coordinates, (425,7), revealing how the model captures the temporal evolution of temperatures in this distinct area. These results showcase the remarkable ability of the DynST-enhanced model to preserve high local fidelity. This fidelity is not just theoretical; it translates into practical, industry-level reliability, consistently maintaining the prediction deviation within a tight 15% margin (Verda et al., 2021). Such performance not only underscores the robustness of the Earthfarseer+DynST model but also highlights its potential for widespread application in scenarios demanding high precision and reliability (Fig 6 also supports our research findings).